# Driving by Hybrid Navigation: An Online HD-SD Map Association Framework and Benchmark for Autonomous Vehicles

### v2

@@ Abstract
Autonomous vehicles rely on global standard-definition (SD) maps for road-level route planning and online local high-definition (HD) maps for lane-level navigation. However, recent work concentrates on construct online HD maps, often overlooking the association of global SD maps with online HD maps for hybrid navigation, making challenges in utilizing online HD maps in the real world. Observing the lack of the capability of autonomous vehicles in navigation, we introduce Online Map Association, the first benchmark for the association of hybrid navigation-oriented online maps, which enhances the planning capabilities of autonomous vehicles. Based on existing datasets, the OMA contains 480k of roads and 260k of lane paths and provides the corresponding metrics to evaluate the performance of the model. Additionally, we propose a novel framework, named Map Association Transformer, as the baseline method, using path-aware attention and spatial attention mechanisms to enable the understanding of geometric and topological correspondences. The code and dataset can be accessed at https://github.com/WallelWan/OMA-MAT.
@@ 1 Introduction
As autonomous driving advances rapidly, the researcher’s focus has been on developing precise and reliable navigation systems. The Standard Definition Map (SD Map) and the High Definition Map (HD Map) served as crucial components in navigation systems, each providing different degrees of environmental detail [9]. The SD Map serves as global navigation at the road level, enjoying widespread application due to its frequent updates and economical storage demands [23], despite its meters-level precision and absence of lane specifics, as illustrated in Fig. 1 (a). On the other hand, HD Map provides detailed geometric information and topology at lane-level to facilitate exact positioning and lane-specific route planning [19], but its extensive production costs and slower update cycles restrict its applicability.
Figure 1: The comparison of navigation on (a) SD map , (b) online HD map and (c) SD-HD map association (Ours). (d) The challenge of map association, including GPS shift, missing and noise of centerline. The identical color in both road and centerline indicates a corresponding pair.
Recently, to address the cost problems of global HD maps, the focus has shifted to online HD map construction [17, 14], aimed at constructing localized HD maps surrounding the vehicle using the vehicle’s own sensing system (refer to Fig. 1 (b)). However, because online HD maps prioritize the view of the ego-vehicle, they are unable to provide a complete centerline topology from start to end, which means they lack the capability for navigation [13]. To enable navigation using the online HD map, it is necessary to convert global navigation at the road-level on the SD map into a local lane-level control strategy corresponding to the online HD map [12], which involves creating an association between the road in the SD map and the lane on the online HD map, as illustrated in Fig. 1 (c). The end result is the successful implementation of hybrid navigation that is based on both the SD map and the online HD map.
However, three significant challenges are encountered during map association [19]. Initially, vehicle positioning error or accuracy issues with SD Map can lead to a GPS shift in location between HD Map and SD Map (as depicted in Fig. 1 (d) (1)), causing location-based KNN matching [5] to frequently produce inaccurate results. Secondly, due to obstructions or disturbances, online HD maps face challenges with intricate topology arising from missing (Fig. 1 (d) (2)) or noise (Fig. 1 (d) (3)) of centerline. Hidden Markov Models (HMM) [25] are ineffective here. Third, a significant challenge in this field is the scarcity of large datasets, which has limited deep learning research in this domain.
To address this fundamental challenge, we thoroughly explore three aspects of contribution: (1) dataset construction, (2) evaluation metric, and (3) algorithm development. Using nuScenes, we introduce the first dataset for the online map association (OMA) with an evaluation metric named Association P-R. OMA derived from nuScenes [3] and OpenStreetMap [1] and boasts more than 30k scenarios and more than 480k roads and 260k lane-level paths. Through a combination of automated and manually calibrated annotations, high-precision road-lane matching annotations are achieved. For the evaluation metrics, we introduce a specific metric called Association P-R, which includes three separate metrics: A-P, A-R, and A-F1. This metric is designed for the association of the online map, taking into account both the accuracy and the precision of the topological alignment.
Furthermore, we present the Map Association Transformer (MAT), a transformer-inspired framework designed to form an online map association utilizing dual attention mechanisms. MAT handles vectorized map inputs by leveraging path-aware attention for learn topological representations and spatial attention to aggregate spatial context. With path-aware and spatial attention, the model is capable of simultaneously learning local geometric correspondences along with global topological structures.
In summary, our contributions are three folds:
  • We introduce Online Map Association (OMA), the first benchmark for hybrid navigation-oriented online map association.
  • We introduce Association P-R, a metric for map association that considers the accuracy and precision of topological alignment.
  • We propose a Map Association Transformer (MAT), which utilizes path-aware attention and spatial attention mechanisms to enable understanding of geometric and topological correspondences.
@@ 2 Related Work
Path Planning. Planning the path for autonomous driving requires coordinated global and local strategies. Global planning identifies optimal routes using graph-based methods such as Dijkstra [8] and A* [10] on SD maps, while local trajectory prediction generates detailed paths through optimal control algorithms [32, 33]. Current datasets [4, 30] focus on local forecasting based on HD maps but lack SD-HD integration. We extend nuScenes [4] with OpenStreetMap SD links and annotate lane-to-link connections, creating the first HD2SD binding dataset.
Online HD Map Construction. The construction of HD maps is a popular topic in autonomous driving and is crucial for subsequent tasks [35, 29, 34]. HDMapNet [14] pioneered BEV-based map generation through sensor fusion, while LSS [26] introduced depth-aware BEV transformation. VectorMapNet [20] enabled end-to-end vector prediction, and the MapTR series [16, 17] introduced hierarchical query embeddings for instance-level construction. We adapt MapTRv2 for online HD construction using our dataset’s annotations and demonstrate its compatibility with SD-HD association.
Map Association. Associating HD lane centerlines to SD road links provides a global context for trajectory prediction. Conventional methods like HMM [25] struggle with lane-level accuracy, while closed PDM [6] lacks explicit map association. We propose a contrastive learning framework that extracts semantic features from HD maps and explicitly associates them with SD links, establishing a robust baseline for cross-map alignment.
@@ 3 Task Definition
Figure 2: (a) The schema of SD map input:Road $\mathcal{R}$. (b) The schema of online HD map input: Centerline $\mathcal{L}$ and boundary $B$. (c) The objective in our task: Mapping function $f$.
In autonomous driving systems, the precise alignment of standard definition (SD) maps with online high definition (HD) maps is critical for lane-level navigation [9, 19]. This task enables planning modules to execute accurate maneuvers by combining real-time HD map observations (e.g., dynamic centerline configurations) with static SD map topologies. We formalize this problem by learning a function that aligns each HD map centerline with its corresponding road in the SD map.
SD Map. As shown in Fig.2 (a), SD maps represent road networks that use roads as primary primitives. Formally, a SD map is defined as a graph $\mathcal{G}_{\mathcal{R}}=(\mathcal{R},\mathcal{E}_{\mathcal{R}})$, where =$\mathcal{R}=\{r_{1},r_{2},\ldots,r_{m}\}$ denotes a set of roads and $\mathcal{E}_{\mathcal{R}}$ encodes their topological connectivity. Each road $r_{j}$ is parameterized by a sequence of directed vectors:
@@
HD Map. As shown in Fig.2 (b), Online HD maps provide details of the lane level, primarily represented as a center line network. We model a HD map as a graph $\mathcal{G}_{\mathcal{L}}=(\mathcal{L},\mathcal{E}_{\mathcal{L}})$, where =$\mathcal{L}=\{l_{1},l_{2},\ldots,l_{n}\}$ is a set of centerlines, each sampled at uniform intervals:
@@
$\mathcal{E}_{\mathcal{L}}$ captures topological relations between the adjacent centerlines. In addition, we include road boundary vectors =$\mathcal{B}=\{b_{1},b_{2},\ldots,b_{m_{b}}\}$, which reflect the extent and shape of the actual road. Each boundary $b_{j}$ is specified as:
@@
Objective. As shown in Fig.2 (c), given $\mathcal{G}_{\mathcal{R}}$ and $\mathcal{G}_{\mathcal{L}}$, the task is to learn a mapping function $f:\mathcal{L}\rightarrow\mathcal{R}$ that assigns each centerline $l\in\mathcal{L}$ to its corresponding road $r_{l}\in\mathcal{R}$. The function satisfies two key constraints:
  1. 1.
    Uniqueness: Each centerline $l$ maps to exactly one ground truth road $r_{l}$;
  2. 2.
    Multiplicity: A single road $r\in\mathcal{R}$ may be associated with multiple centerlines $l_{1},l_{2},\ldots\in\mathcal{L}$.
This formulation casts the alignment task as a many-to-one classification problem, where the number of classes equals the number of $|\mathcal{R}|$, and each centerline acts as an input sample. The goal is to maximize classification accuracy while preserving topological consistency between HD and SD maps.
@@ 4 Dataset and Metric
In this section, we provide a summary of the Online Map Association (OMA) dataset. The Online Map Association (OMA) dataset is built on nuScenes [4] and OpenStreetMap [1] (CC BY-NC-SA 4.0), containing 30K+ HD-SD map pairs with high-quality annotations generated via automated labeling, manual refinement, and multi-stage validation. Dataset statistics are summarized in supplemental material.
Figure 3: (a) The visualization of SD map and HD map with association annotations of Boston in nuScenes. The same color implies an associative pair. (b) The visualization of OMA. (c) The visualization of OMA-GT.
4.1 Dataset Construction
Raw Data and Annotation. The source of raw HD maps is nuScenes [4], which includes locations in Boston and Singapore, featuring centerline geometries scanned with LiDAR. For these areas, SD maps were obtained from OpenStreetMap (OSM). Initially, GPS coordinates are used to align HD and SD maps, after which manual adjustments are made to address any remaining misalignments due to GPS inaccuracies or differences in map projections. The visualization is shown in Fig. 3 (a), which shows the global SD and HD map of Boston after map association.
An automated labeling pipeline using a Map Association Transformer (MAT, Section 5) trained on proprietary Chinese HD-SD maps generates initial annotations on nuScenes, followed by topological post-processing. Professional annotators refine these drafts through geometry and topology adjustments under domain-specific rules. The result is high-quality annotations that meet real-world deployment needs.
OMA. To evaluate robustness to predicted HD maps, the online HD map in OMA is generated by MapTRv2 [17], as shown in Fig 3 (b). The model is trained on nuScenes using synchronized LiDAR and camera inputs with offical configuration. For each sample:
  • SD Map: An SD map cropping measure $150\,m\times 150\,m$ centered around the ego vehicle preserves the adjacent topological context.
  • HD Map: An HD map cropping measure $30\,m\times 60\,m$ centered around the ego vehicle, as referenced by [17, 21, 15], is used as input for the online HD map.
OMA-GT. We propose OMA-GT to further assess the robustness of association algorithms against map dependencies. Unlike OMA, OMA-GT is designed by emulating online perception through the use of localized ground truth HD maps surrounding the ego vehicle, as depicted in Fig 3 (c).
Figure 4: Example of TP, FP and FN for evaluate Association Precision-Recall.
@@ 4.2 Evaluation Metric
Existing lane-level accuracy metrics do not reflect global navigation performance. To address this, we propose Association Precision-Recall (Association P-R) for a comprehensive evaluation of the SD-HD map Association quality.
Reachability P-R [22]: This metric evaluates the connectivity of the path between landmarks. A predicted path between locations $\hat{A}$ and $\hat{B}$ is considered a true positive (TP) if its Chamfer distance to any ground-truth path between corresponding $A$-$B$ pairs is under threshold, regardless of direct connectivity between $\hat{A}$-$\hat{B}$.
Association P-R: The Association P-R is introduced with two significant upgrades based on Reachability P-R, as illustrated in Fig. 4:
  1. 1.
    Label Sequence Alignment: Both predicted and ground-truth paths are converted to simplified label sequences representing SD map link traversals.
  2. 2.
    Length-aware Overlap Check: For each aligned label, we calculate the overlap ratio between the predicted and ground-truth path segments. An TP is confirmed when this ratio exceeds the threshold $T$.
Following the mAP conventions [18], we use $T=[0.5:0.05:0.95]$ (10 thresholds) and report mean P-R and F1 scores. To mitigate path-length bias, we separately compute metrics across 15 length intervals $L=[[0,5),[5,10),..,[75,+\infty)]$ before aggregation. Specifically, in OMA-GT, metrics such as A-R and A-F1 become irrelevant because M-R achieves 100% due to the same HD map as the ground truth. The detailed of Association P-R are provided in supplemental material.

### v3

# Online Navigation Refinement: Achieving Lane-Level Guidance by Associating Standard-Definition and Online Perception Maps

@@ Abstract
Lane-level navigation is critical for geographic information systems and navigation-based tasks, offering finer-grained guidance than road-level navigation by standard definition (SD) maps. However, it currently relies on expansive global HD maps that cannot adapt to dynamic road conditions. Recently, online perception (OP) maps have become research hotspots, providing real-time geometry as an alternative, but lack the global topology needed for navigation. To address these issues, Online Navigation Refinement (ONR), a new mission is introduced that refines SD-map-based road-level routes into accurate lane-level navigation by associating SD maps with OP maps. The map-to-map association to handle many-to-one lane-to-road mappings under two key challenges: (1) no public dataset provides lane-to-road correspondences; (2) severe misalignment from spatial fluctuations, semantic disparities, and OP map noise invalidates traditional map matching. For these challenges, We contribute: (1) Online map association dataset (OMA), the first ONR benchmark with 30K scenarios and 2.6M annotated lane vectors; (2) MAT, a transformer with path-aware attention to aligns topology despite spatial fluctuations and semantic disparities and spatial attention for integrates noisy OP features via global context; and (3) NR P-R, a metric evaluating geometric and semantic alignment. Experiments show that MAT outperforms existing methods at 34 ms latency, enabling low-cost and up-to-date lane-level navigation.
Figure 1: Motivation for online navigation refinement and map association. The roads and lanes with the same color indicate that they are interconnected: (a) Standard Definition (SD) maps offer only road-level navigation without lane details. (b) Online Perception (OP) maps offer lane-specific details, yet they are not connected to SD maps and cannot identify the correct lane. (c) Associating SD and OP maps enables lane selection, achieving online navigation refinement and lane-level navigation.
[…]
@@ 1 Introduction
Lane-level navigation has emerged as a critical capability in geographic information systems (GIS) hansson2020lane; guo2025lane and navigation-based tasks peng2025navigscene, offering finer-grained guidance than traditional road-level navigation based on standard definition (SD) maps (Fig. 1 (a)) zhang2024survey. However, today’s lane-level navigation systems mostly rely on pre-built HD maps, which are expensive to create and maintain. They often miss real-world changes elghazaly2023high—like construction, rule updates, or accidents—leading to outdated guidance and safety risks. Recently, online perception (OP) maps have become research hotspots within autonomous driving li2022hdmapnet; maptrv2. Powered by vehicle sensors, the OP maps offer current and localized lane depictions. However, as shown in Fig. 1 (b), they inherently lack global route topology and cannot support lane-level navigation wong2020mapping.
To achieve up-to-date and low-cost lane-level navigation, we introduce a new mission called Online Navigation Refinement (ONR). The core objective of the mission is to transform the road-level navigation derived from SD maps into precise lane-level navigation aligned with the OP maps. For a fast and accurate process, ONR needs a paradigm of map-to-map matching, called map association, as shown in Fig 1 (c). Compared with map matching ($M$) that bind GPS trajectories to static SD/HD maps chao2020survey as a path-to-map matching based on the HMM DBLP:conf/gis/NewsonK09, seq-to-seq model feng2020deepmm; ren2021mtrajrec, graph model liu2023graphmm and transformers tang2025elevation, map association uses map-to-map matching because SD maps and OP maps are heterogeneous and do not share a one-to-one correspondence. Attempting to directly align SD paths with OP maps, or the other way around, overlooks the semantic disparities between the two.
Specifically, map association faces two key challenges. First, no public datasets provide structured, lane-to-road-level correspondences between SD maps and online perception (OP) maps. Although there exist auto-driving datasets (e.g., nuScenes caesar2020nuscenes, OpenLaneV2 wang2023openlanev2) offer local lane geometries, and OpenStreetMap provides road-level topologies, none establish explicit, learnable mappings between them. Secondly, SD and OP maps display inherent heterogeneity due to diverse granularity, resulting in substantial spatial and semantic differences. A robust association mechanism is required to manage spatial fluctuations (such as GPS drift and changes in scale) and semantic disparities (such as the differing number of lanes on a road). It should also be capable of processing noise issues in OP maps in real time, which include lane discontinuities, omissions, or errors in OP mapping. Figure 6 in Appendix D illustrates the complexity and challenge of map association.
To address these challenges, we make three core contributions: (1) Dataset: We introduce Online Map Association Dataset (OMA), the first open source benchmark for the online navigation refinement mission in the paradigm of map association, derived from nuScenes caesar2020nuscenes and OpenStreetMap osm, which contains more than 30K scenarios, 480K road paths and 2.6M lane vectors with manually annotated associations. (2) Baseline: We present MAT, a lightweight transformer-based model for the real-time association of maps. To handle spatial and semantic differences, MAT incorporates two key modules: path-aware attention and spatial attention. Path-aware attention reorders and group vector tokens by path index, implicitly encoding topological structure to align SD and OP maps despite spatial offsets. Spatial attention sorts and groups road/centerline tokens using a spatial curve, enabling global context modeling through cross-topological feature integration. (3) Metric: We propose Navigation Refinement P-R (NR P-R), a metric that measures alignment of the path and the correspondence through precision, recall and F1 specifically designed for map association. Traditional $M$ metrics only measure the accuracy of the matching, but the NR P-R evaluates both the geometric similarity and the accuracy of the matching. Moreover, NR P-R only utilizes annotation in ground-truth perception maps, rendering the benchmark suitable for the evaluation of any map generation method.
Extensive experiments indicate that MAT achieves state-of-the-art performance with 34ms latency on OMA, showing significant improvements compared to traditional map matching methods and deep learning map matching method, enabling low-cost and up-to-date lane-level navigation.
@@ 2 Related Work
Map Matching. Map matching investigates how to link GPS tracks to SD or HD maps. hansson2020lane utilize a hidden Markov model (HMM) considering trajectories as observations and the corresponding road segments as states, demonstrating superiority over earlier geometric matching algorithms. In the realm of deep learning, the main approach involves sequence-to-sequence (Seq2Seq) methods feng2020deepmm; ren2021mtrajrec. With an input trajectory consisting of a sequence of geo-points, the encoder-decoder framework provides a sequence of matched road segments. Recently, GraphMM liu2023graphmm has used a graph-based approach that explicitly integrates all the correlations mentioned above, while EAM3 tang2025elevation use a BERT-like transformer alaparthi2020bidirectional to achieve the best performance. Compared to the map matching method, MAT uses a map-to-map paradigm to accommodate the many-to-one correspondence between lanes and roads in real time. Moreover, unlike the self-attention used in EAM3, MAT employs spatial and path-aware attention by incorporating the spatial curve and the path index, thus striking a balance between performance and latency.
Map Generation and Priors. The construction of online perception maps is a popular topic in autonomous driving hao2024mapdistill; zhang2024sparse; li2024mlp and is crucial for subsequent tasks wan2024sp; zhang2024p2ftrack. HDMapNet li2022hdmapnet pioneered BEV-based map generation through sensor fusion, while LSS philion2020lift introduced depth-aware BEV transformation. VectorMapNet liu2023vectormapnet enabled end-to-end vector prediction, and the MapTR liao2023maptr series introduced hierarchical query embeddings for instance-level construction. We adapt MapTRv2 maptrv2 for online HD construction using the annotations of our dataset and demonstrate its compatibility with the SD-HD association. Recent works such as SMERF luo2024augmenting, P-MapNet jiang2024p, and TopoSD yang2024toposd leverage SD map priors to mitigate sensor noise and refine perception geometry. Unlike these approaches that focus on enhancing generation quality, MAT targets the explicit association phase, utilizing topological optimization to resolve assignment ambiguity arising from residual noise.
@@ 3 Task Definition
Figure 2: (a) The schema of SD map input: Road $\mathcal{R}$. (b) The schema of OP map input: Centerline $\mathcal{L}$ and boundary $B$. (c) The objective in our task: Mapping function $f$.
The goal of online navigation refinement is to transform road-topology-aligned navigation into lane-topology-aligned navigation. To achieve this, we utilize the map association as a map-to-map match paradigm. This involves correlating SD maps with online perception maps, allowing us to translate any route from an SD map into the corresponding route on an online perception map.
Standard Definition Map (SD Map). As shown in Fig.2 (a), SD maps represent road networks that use roads as primary primitives. Formally, a SD map is defined as a graph $\mathcal{G}_{\mathcal{R}}=(\mathcal{R},\mathcal{E}_{\mathcal{R}})$, where $\mathcal{R}=\{r_{1},r_{2},\ldots,r_{m}\}$ denotes a set of roads and $\mathcal{E}_{\mathcal{R}}$ encodes their topological connectivity. Each road $r_{j}$ is parameterized by a sequence of directed vectors:
@@
Online Perception Map (OP Map). As shown in Fig.2 (b), OP maps provide details of the lane level, primarily represented as a center line network. We model an OP map as a graph $\mathcal{G}_{\mathcal{L}}=(\mathcal{L},\mathcal{E}_{\mathcal{L}})$, where $\mathcal{L}=\{l_{1},l_{2},\ldots,l_{n}\}$ is a set of centerlines, each sampled at uniform intervals:
@@
$\mathcal{E}_{\mathcal{L}}$ captures topological relations between the adjacent centerlines. In addition, we include road boundary vectors $\mathcal{B}=\{b_{1},b_{2},\ldots,b_{m_{b}}\}$, which reflect the extent and shape of the actual road. Each boundary $b_{j}$ is specified as:
@@
Objective. As shown in Fig.2 (c), given $\mathcal{G}_{\mathcal{R}}$ and $\mathcal{G}_{\mathcal{L}}$, the task is to learn a mapping function $f:\mathcal{L}\rightarrow\mathcal{R}$ that assigns each centerline $l\in\mathcal{L}$ to its corresponding road $r_{l}\in\mathcal{R}$. The function satisfies two key constraints: 1. Uniqueness: Each centerline $l$ corresponds to exactly one ground truth road $r_{l}$; 2. Multiplicity: A single road $r\in\mathcal{R}$ may be associated with multiple centerlines $l_{1},l_{2},\ldots\in\mathcal{L}$.
This formulation casts the alignment task as a multi-to-one classification problem, where the number of classes equals the number of $|\mathcal{R}|$, and each centerline acts as an input sample. The goal is to maximize classification accuracy while preserving topological consistency between SD and OP maps. After we form this association, any path on the SD map allows us to locate all matching paths on the OP map through a single topological sorting.
@@ 4 Dataset and Metric
Figure 3: (a) The visualization of SD map and GT OP map with association annotations of Boston in nuScenes. The same color implies an associative pair. (b) The visualization of OMA Train/Val set. (c) The visualization of OMA Test set.
4.1 Dataset Overview
The source of ground-truth OP maps is nuScenes nuscenes2019, which includes locations in Boston and Singapore, featuring centerline geometries scanned with LiDAR. The SD maps were obtained from OpenStreetMap (OSM). We manually annotated the data to establish the association between the ground truth of the OP map and the SD map. The visualization is shown in Fig. 3 (a), which shows the global SD and ground-truth OP map of Boston with an association annotation.
The dataset is divided into training, validation, and test sets (see Fig. 3 (b) and (c)). Within the nuScenes framework, we employ the pon-split of nuScene. This approach designates distinct areas for training and validation/testing datasets, effectively preventing data leakage. The validation and test data sets are generated from the same scenes, but they have a variety of OP maps. For the OP map, both training and validation use ground-truth maps with manual correspondence annotations. For testing, the OP map is generated by MapTRv2 and its annotation is the same as the scene in the val set. It is essential to explain that the validation set and test set in OMA have equal significance. The purpose of the validation set is to assess the association capability of the OP map under noise-free conditions, whereas the test set evaluates this capability when the OP map experiences significant noise. This distinction sets apart the validation/test split in OMA from that in typical datasets. The specific analysis is provided in the Appendix  D.
Figure 4: Example of TP, FP and FN for evaluate Navigation Refinement Precision-Recall.
@@ 4.2 Evaluation Metric
In map matching, evaluation typically involves metrics such as Trajectory-Level Accuracy and the Ratio of Longest Common Subsequence, which presume a one-to-one mapping between trajectories and paths, concentrating only on the accuracy and sequence of path matching. Conversely, in map association, to assess various map generation methods, our goal is to use the association annotations of the ground-truth OP map to evaluate the results of the various OP maps with noise. Drawing inspiration from Reachability Precision Recall (P-R) lu2023translating, we developed the Navigation Refinement P-R metric. This allows for the assessment of any map generation method’s perception map performance for the map association task, relying solely on ground-truth OP map annotations.
Navigation Refinement P-R: Fig. 4 provides a detailed description of typical scenarios encountered during the Navigation Refinement Precision Recall (P-R) evaluation. The examination protocol involves three consecutive stages:
1. Path Matching between Prediction and Ground Truth: In the first stage, a connection is drawn between the predicted paths and GT paths within the OP map. Notably, only GT paths come with association annotations, making this matching essential for the following evaluations. A match is deemed successful if the bidirectional Chamfer distance is less than the distance threshold of 1 meter, consistent with Reachability P-R lu2023translating. Predictions that do not meet this threshold are labeled FP, thereby excluding them from further SD-HD analysis. As illustrated in Row 1 of Fig. 4, Case 1 (left) shows that the prediction with minimal deviation successfully aligns with the GT paths. While Case 1 (middle) represents the heavily-deviated prediction results in an FP label.
2. Label Sequence Alignment Verification: Successful matches are converted into simplified label sequences, indicating SD map link traversals. Failures in sequence alignment at this point lead to FP/FN labels. As shown in Case 2 (middle) of Fig. 4, the case is a misaligned sequence result which is an FP, while Case 2 (right) shows that the GT without a match leads to a false negative (FN).
3. Length-Aware Overlap Assessment: For aligned sequences, we quantify the semantic consistency by computing the Label Overlap score. Specifically, this score is defined as the ratio of the length of correctly matched segments (where the predicted road ID matches the ground truth road ID) to the total path length. We compare the length proportions of road labels between the predicted and GT sequences to determine these matches. The true positive (TP) classification is determined based on thresholds. As shown in Case 3 (left) of Fig. 4, the case is a 67% overlap, which is TP at T = 50% but FP at T=95%.
Following the mAP conventions lin2014microsoft, we use $T=[0.5:0.05:0.95]$ (10 thresholds) and report mean P-R and F1 scores. To mitigate path-length bias, we separately compute metrics across 15 length intervals $L=[[0,5),[5,10),..,[70,+\infty)]$ before aggregation. The ablation study of distance threshold and length intervals is shown in the appendix G.